# PairFlow: Closed-Form Source-Target Coupling for Few-Step Generation in Discrete Flow Models

### v1

@@ E.2 Continuous flow matching on MNIST
We train rectified flow models on MNIST (lecun2002gradient) using two pairing strategies: (i) independent pairing (baseline) and (ii) closed-form pairing as described in Sec. E.1. We adopt CondOT (Lipman:2023FlowMatching) as our base flow model, which is originally trained with a independent pairing. We denote the variant of CondOT trained on pairs generated by the closed-form forward velocity as PairFlow. To enable a few-step sampling, we subsequently apply rectification distillation (ReFlow (Liu:2023RF)) to each pretrained model, denoted by the suffix “+RF”.
We use an NCSN++-style U-Net backbone (song2021scorebased) with a base width of 64 and 3 downsampling stages (doubling channels at each stage), optimized using Adam (Kingma:2014Adam) with a learning rate of $2\times 10^{-4}$. The pretraining takes 500 epochs. The distillation stage requires 200 epochs with a learning rate of $2\times 10^{-5}$.
Tab. 14 summarizes performance at various sampling steps. Without distillation, closed-form pairing (PairFlow) yields significantly better FID in the few-step settings and maintains the performance in the many-step settings, relative to the baseline. With distillation (ReFlow (Liu:2023RF), our method still shows better performance: PairFlow+RF achieves a lower FID in every sampling budget than ReFlow applied to the baseline. These results show that closed-form pairing benefits both undistilled and distilled flow models, with especially large gains when the sampling steps are small.
@@ Appendix F Additional Qualitative Results
In addition to Fig. 4 in the main paper, we further visualize the 1-step and 2-step generation results for MNIST-Binary (lecun2002gradient) in Fig. 9 and Fig. 10. As discussed in Sec. 5.3PairFlow outperforms the base models (Schiff:2025UDLM; Sahoo:2024MDLM) and achieves comparable quality to the base models combined with acceleration methods (Sahoo:2025Duo; Yoo:2025ReDi). Additional visualizations for CIFAR-10 (krizhevsky2009learning) with 64- and 256-step generations are shown in Fig. 8, demonstrating that our method outperforms the other base models.
@@ Appendix G Re-Flow Iteration results
This section details the results of applying the iterative re-flow procedure (liuflow; Yoo:2025ReDi) on the QM9 dataset (ramakrishnan2014quantum). We generated 1,024 samples across various timesteps and re-flow iterations, assessing validity, uniqueness, and novelty. The configuration for subsequent re-flow iterations follows the same protocol as our main experiment. As shown in Tab. 15, Tab. 16, and Tab. 17, all metrics are averaged over 10 independent runs with standard deviations provided. We observed that iterative re-flow consistently enhances few-step generation capabilities. Notably, PairFlow demonstrates superior performance over the baseline; it outperforms UDLM not only at equivalent iteration levels but also surpasses UDLM with multiple re-flow iterations even when PairFlow uses one or no additional iterations.
Table 15: validity scores ($\uparrow$) on QM9 (ramakrishnan2014quantum) for UDLM and PairFlow across varying rectification steps and NFEs (1 to 64). The best and second-best values per column are highlighted in bold and underlined, respectively.
@@
Table 16: Uniqueness scores ($\uparrow$) on QM9 (ramakrishnan2014quantum) for UDLM and PairFlow across varying rectification steps and NFEs (1 to 64). The best and second-best values per column are highlighted in bold and underlined, respectively.
@@
Table 17: Novelty scores ($\uparrow$) on QM9 (ramakrishnan2014quantum) for UDLM and PairFlow across varying rectification steps and NFEs (1 to 64). The best and second-best values per column are highlighted in bold and underlined, respectively.
@@
In this section, we present comprehensive experimental results from applying our subset-partition pairing technique to the ZINC-250k molecular dataset (irwin2012zinc). Following the same protocol as our main experiments, we generated 1,024 samples across varying timesteps and partition counts to evaluate validity, uniqueness, and novelty. The only deviation from the standard method is the pairing strategy; here, we calculate the closed-form backward velocity exclusively within each subset to reduce the computational cost of Eqn. 10. All reported metrics, summarized in Tab. 18, Tab. 19, and Tab. 20, are averaged over 10 independent runs with corresponding standard deviations. Additionally, we report the pairing-time cost for each subset-partition configuration. The results demonstrate that our subset-pairing algorithm effectively reduces the computational time for pairing, while maintaining performance comparable to the full-set baseline.
Table 18: Validity scores ($\uparrow$) on the Zinc-250k dataset (irwin2012zinc) evaluated across different subset partitions and NFEs (1 to 64). The best and second-best values per column are highlighted in bold and underlined, respectively. $T_{\textsc{PairFlow}{}}$ denotes the runtime (in minutes) of each configuration, measured in wall-clock time using an RTX A6000 GPU.
@@
Table 19: Uniqueness scores ($\uparrow$) on Zinc-250k (irwin2012zinc) evaluated across different subset partitions and NFEs (1 to 64). The best and second-best values per column are highlighted in bold and underlined, respectively. $T_{\textsc{PairFlow}{}}$ denotes the runtime (in minutes) of each configuration.
@@
Table 20: Novelty scores ($\uparrow$) on Zinc-250k (irwin2012zinc) evaluated across different subset partitions and NFEs (1 to 64). The best and second-best values per column are highlighted in bold and underlined, respectively. $T_{\textsc{PairFlow}{}}$ denotes the runtime (in minutes) of each configuration.
@@ Appendix I Application for more Complex Systems
In this section, we evaluate our method on a higher-dimensional dataset. Specifically, we use the FFHQ (karras2019styleffhq) dataset, downsampled to $64\times 64$. Following the same protocol as in our main experiments, we generate 5,000 samples across varying timesteps and report the FID computed against the training set. The results of this experiment are provided in Tab. 21. All training hyperparameters are kept identical to those used in the CIFAR-10 experiments described in Section 5.
Table 21: Comparison of FID scores ($\downarrow$) on FFHQ (karras2019styleffhq) downsampled to $64\times 64$ resolution across extended NFE steps (1 to 1024). Best values per column are highlighted in bold.
@@
We adopt the LM1B (chelba2013exploringlm1b) dataset to evaluate our method under a substantially larger vocabulary size and training corpus. The text corpus is segmented into sequences of varying lengths ($N=16,32,64,128$), while keeping the total number of training samples fixed ($|X_{1}|\approx 3.5\text{M}$). To assess generation quality, we compute generative perplexity using GPT-2 Large and entropy on 1,024 generated samples for each NFE setting. The results are summarized in Tab. 22 and Tab. 23. For training, we follow the network hyperparameter configuration of (Schiff:2025UDLM), modifying only the number of training iterations for each sequence dimensionality.
Table 22: Generative Perplexity ($\downarrow$) on LM1B (chelba2013exploringlm1b) measured with GPT2-large across varying lengths (NN) and their corresponding training iterations (Iter.) over NFE steps 4 to 1024. Best values are highlighted in bold.
@@
Table 23: Comparison of Entropy ($\uparrow$) on LM1B (chelba2013exploringlm1b) across varying lengths (NN) and training iterations (Iter.) over NFE steps 4 to 1024. Best values are highlighted in bold.
@@
In this section, we evaluate our method using FID computed on the test sets of two image domains: CIFAR-10 (krizhevsky2009learning) and MNIST-Binary (lecun2002gradient). For CIFAR-10, we additionally report FID scores measured with DINOv2 (oquab2024dinov2). The overall results are summarized in Tab. 24 and Tab. 25. Across all evaluation metrics, the performance trend is consistent with our main findings—PairFlow delivers improved generation quality over the baseline, with especially strong gains in the few-step generation regime.
Table 24: FID and FID-Dino scores ($\downarrow$) on test dataset for CIFAR-10 (krizhevsky2009learning) comparison across extended NFE steps (1 to 1024). Best values are bolded.
@@
Table 25: FID scores ($\downarrow$) on the MNIST-Binary (lecun2002gradient) test set across various NFE steps (1 to 64). Best values are bolded.
@@
We further assess potential memorization by measuring nearest-neighbor distances with respect to the training set. For MNIST-Binary (lecun2002gradient), we compute pixel-wise $\ell_{2}$ distances, whereas for CIFAR-10 (krizhevsky2009learning), we evaluate both $\ell_{2}$ distance and cosine similarity between features extracted using DINOv2 (oquab2024dinov2). As summarized in Tab. 26 and Tab. 27, across all evaluation settings, the nearest-neighbor distances of PairFlow are comparable to or slightly larger than those of the baseline. These results support the conclusion that our method does not suffer from severe overfitting or excessive memorization of the training data.
Table 26: Comparison of $\ell_{2}$ and Dino (oquab2024dinov2) Cosine nearest neighbor distance on the CIFAR-10 (krizhevsky2009learning) training set across extended NFE steps (1 to 1024). Best values are bolded.
@@
Table 27: Comparison of $\ell_{2}$ nearest neighbor distance on the MNIST-Binary (lecun2002gradient) training set across extended NFE steps (1 to 64). Best values are bolded.

### v2

@@ E.2 Continuous flow matching on MNIST
We train rectified flow models on MNIST (LeCun et al., 2002) using two pairing strategies: (i) independent pairing (baseline) and (ii) closed-form pairing as described in Sec. E.1. We adopt CondOT (Lipman et al., 2023) as our base flow model, which is originally trained with a independent pairing. We denote the variant of CondOT trained on pairs generated by the closed-form forward velocity as PairFlow. To enable a few-step sampling, we subsequently apply rectification distillation (ReFlow (Liu et al., 2023)) to each pretrained model, denoted by the suffix “+RF”.
We use an NCSN++-style U-Net backbone (Song et al., 2021) with a base width of 64 and 3 downsampling stages (doubling channels at each stage), optimized using Adam (Kingma and Ba, 2014) with a learning rate of $2\times 10^{-4}$. The pretraining takes 500 epochs. The distillation stage requires 200 epochs with a learning rate of $2\times 10^{-5}$.
Tab. 14 summarizes performance at various sampling steps. Without distillation, closed-form pairing (PairFlow) yields significantly better FID in the few-step settings and maintains the performance in the many-step settings, relative to the baseline. With distillation (ReFlow (Liu et al., 2023), our method still shows better performance: PairFlow+RF achieves a lower FID in every sampling budget than ReFlow applied to the baseline. These results show that closed-form pairing benefits both undistilled and distilled flow models, with especially large gains when the sampling steps are small.
@@ Appendix F Additional Qualitative Results
In addition to Fig. 4 in the main paper, we further visualize the 1-step and 2-step generation results for MNIST-Binary (LeCun et al., 2002) in Fig. 9 and Fig. 10. As discussed in Sec. 5.3PairFlow outperforms the base models (Schiff et al., 2025; Sahoo et al., 2024) and achieves comparable quality to the base models combined with acceleration methods (Sahoo et al., 2025; Yoo et al., 2025). Additional visualizations for CIFAR-10 (Krizhevsky and others, 2009) with 64- and 256-step generations are shown in Fig. 8, demonstrating that our method outperforms the other base models.
@@ Appendix G Re-Flow Iteration results
This section details the results of applying the iterative re-flow procedure (Liu et al., 2023; Yoo et al., 2025) on the QM9 dataset (Ramakrishnan et al., 2014). We generated 1,024 samples across various timesteps and re-flow iterations, assessing validity, uniqueness, and novelty. The configuration for subsequent re-flow iterations follows the same protocol as our main experiment. As shown in Tab. 15, Tab. 16, and Tab. 17, all metrics are averaged over 10 independent runs with standard deviations provided. We observed that iterative re-flow consistently enhances few-step generation capabilities. Notably, PairFlow demonstrates superior performance over the baseline; it outperforms UDLM not only at equivalent iteration levels but also surpasses UDLM with multiple re-flow iterations even when PairFlow uses one or no additional iterations.
Table 15: validity scores ($\uparrow$) on QM9 (Ramakrishnan et al., 2014) for UDLM and PairFlow across varying rectification steps and NFEs (1 to 64). The best and second-best values per column are highlighted in bold and underlined, respectively.
@@
Table 16: Uniqueness scores ($\uparrow$) on QM9 (Ramakrishnan et al., 2014) for UDLM and PairFlow across varying rectification steps and NFEs (1 to 64). The best and second-best values per column are highlighted in bold and underlined, respectively.
@@
Table 17: Novelty scores ($\uparrow$) on QM9 (Ramakrishnan et al., 2014) for UDLM and PairFlow across varying rectification steps and NFEs (1 to 64). The best and second-best values per column are highlighted in bold and underlined, respectively.
@@
In this section, we present comprehensive experimental results from applying our subset-partition pairing technique to the ZINC-250k molecular dataset (Irwin et al., 2012). Following the same protocol as our main experiments, we generated 1,024 samples across varying timesteps and partition counts to evaluate validity, uniqueness, and novelty. The only deviation from the standard method is the pairing strategy; here, we calculate the closed-form backward velocity exclusively within each subset to reduce the computational cost of Eqn. 10. All reported metrics, summarized in Tab. 18, Tab. 19, and Tab. 20, are averaged over 10 independent runs with corresponding standard deviations. Additionally, we report the pairing-time cost for each subset-partition configuration. The results demonstrate that our subset-pairing algorithm effectively reduces the computational time for pairing, while maintaining performance comparable to the full-set baseline.
Table 18: Validity scores ($\uparrow$) on the Zinc-250k dataset (Irwin et al., 2012) evaluated across different subset partitions and NFEs (1 to 64). The best and second-best values per column are highlighted in bold and underlined, respectively. $T_{\textsc{PairFlow}{}}$ denotes the runtime (in minutes) of each configuration, measured in wall-clock time using an RTX A6000 GPU.
@@
Table 19: Uniqueness scores ($\uparrow$) on Zinc-250k (Irwin et al., 2012) evaluated across different subset partitions and NFEs (1 to 64). The best and second-best values per column are highlighted in bold and underlined, respectively. $T_{\textsc{PairFlow}{}}$ denotes the runtime (in minutes) of each configuration.
@@
Table 20: Novelty scores ($\uparrow$) on Zinc-250k (Irwin et al., 2012) evaluated across different subset partitions and NFEs (1 to 64). The best and second-best values per column are highlighted in bold and underlined, respectively. $T_{\textsc{PairFlow}{}}$ denotes the runtime (in minutes) of each configuration.
@@ Appendix I Application for more Complex Systems
In this section, we evaluate our method on a higher-dimensional dataset. Specifically, we use the FFHQ (Karras et al., 2019) dataset, downsampled to $64\times 64$. Following the same protocol as in our main experiments, we generate 5,000 samples across varying timesteps and report the FID computed against the training set. The results of this experiment are provided in Tab. 21. All training hyperparameters are kept identical to those used in the CIFAR-10 experiments described in Section 5.
Table 21: Comparison of FID scores ($\downarrow$) on FFHQ (Karras et al., 2019) downsampled to $64\times 64$ resolution across extended NFE steps (1 to 1024). Best values per column are highlighted in bold.
@@
We adopt the LM1B (Chelba et al., 2013) dataset to evaluate our method under a substantially larger vocabulary size and training corpus. The text corpus is segmented into sequences of varying lengths ($N=16,32,64,128$), while keeping the total number of training samples fixed ($|X_{1}|\approx 3.5\text{M}$). To assess generation quality, we compute generative perplexity using GPT-2 Large and entropy on 1,024 generated samples for each NFE setting. The results are summarized in Tab. 22 and Tab. 23. For training, we follow the network hyperparameter configuration of (Schiff et al., 2025), modifying only the number of training iterations for each sequence dimensionality.
Table 22: Generative Perplexity ($\downarrow$) on LM1B (Chelba et al., 2013) measured with GPT2-large across varying lengths (NN) and their corresponding training iterations (Iter.) over NFE steps 4 to 1024. Best values are highlighted in bold.
@@
Table 23: Comparison of Entropy ($\uparrow$) on LM1B (Chelba et al., 2013) across varying lengths (NN) and training iterations (Iter.) over NFE steps 4 to 1024. Best values are highlighted in bold.
@@
In this section, we evaluate our method using FID computed on the test sets of two image domains: CIFAR-10 (Krizhevsky and others, 2009) and MNIST-Binary (LeCun et al., 2002). For CIFAR-10, we additionally report FID scores measured with DINOv2 (Oquab et al., 2024). The overall results are summarized in Tab. 24 and Tab. 25. Across all evaluation metrics, the performance trend is consistent with our main findings—PairFlow delivers improved generation quality over the baseline, with especially strong gains in the few-step generation regime.
Table 24: FID and FID-Dino scores ($\downarrow$) on test dataset for CIFAR-10 (Krizhevsky and others, 2009) comparison across extended NFE steps (1 to 1024). Best values are bolded.
@@
Table 25: FID scores ($\downarrow$) on the MNIST-Binary (LeCun et al., 2002) test set across various NFE steps (1 to 64). Best values are bolded.
@@
We further assess potential memorization by measuring nearest-neighbor distances with respect to the training set. For MNIST-Binary (LeCun et al., 2002), we compute pixel-wise $\ell_{2}$ distances, whereas for CIFAR-10 (Krizhevsky and others, 2009), we evaluate both $\ell_{2}$ distance and cosine similarity between features extracted using DINOv2 (Oquab et al., 2024). As summarized in Tab. 26 and Tab. 27, across all evaluation settings, the nearest-neighbor distances of PairFlow are comparable to or slightly larger than those of the baseline. These results support the conclusion that our method does not suffer from severe overfitting or excessive memorization of the training data.
Table 26: Comparison of $\ell_{2}$ and Dino (Oquab et al., 2024) Cosine nearest neighbor distance on the CIFAR-10 (Krizhevsky and others, 2009) training set across extended NFE steps (1 to 1024). Best values are bolded.
@@
Table 27: Comparison of $\ell_{2}$ nearest neighbor distance on the MNIST-Binary (LeCun et al., 2002) training set across extended NFE steps (1 to 64). Best values are bolded.